Question 19 A country decides to specialize in wheat production. They now export wheat and import shirts. Which groups of people in this country benefit from trade? PICK all that are correct.

A) wheat consumers

B) wheat producers

C) shirt consumers

D) shirt producers

Question 20 When countries specialize in producing certain goods and then freely exchange those goods for other goods with different countries, what is the advantage?

A) Each country can consume at a point outside their production possibilities frontier.

B) All people in each country will be better off than they would be if they did not trade.

C) Each country can produce and consume at a point outside their production possibilities frontier.

D) Each country can produce at a point outside their production possibilities frontier.

Question 21 Susan can brew 5 gallons of root beer in an hour or she can make 4 pizzas in an hour. Becky can brew 7 gallons of root beer in an hour or she can make 5 pizzas in an hour. Who has a comparative advantage in making root beer? Becky or Susan

Question 24 Consider the following production possibilities table that shows different combinations of two goods that can be produced with given resources:

Shirts Corn
A 0 20
B 10 15
C 20 10
D 30 5
E 40 0

When the country moves from point A to point B, the opportunity cost of producing 1 shirt is ______ bushels of corn.

When the country moves from point C to point D, the opportunity cost of producing 1 shirt is ______ bushels of corn. Enter numbers rounded to two decimals as necessary.

Question 22 Two countries produce two goods and they each have 4 million hours of labor. Consider the table that shows the quantity produced per labor hour:

Shirts Corn
Alpha 4 12
Beta 2 10

If each country allocates 2 million hours of labor to each good then in total the two countries can together produce __ million shirts and ___ million bushels of corn. (Enter whole numbers)
Answer #1

19) A, B and C

Consumers of both shirt and wheat will benefit because of the increase in consumption possibilities after specialisation. Producers of wheat will also benefit. Producers of shirt will be worse off as they cannot produce shirts now.

20) A is correct

After specialisation consumption possibilities increase which means people can consume outside their production possibility level.

21) Becky

Susan's opportunity cost of making root beer = 4/5= 0.8 pizzas

Becky's opportunity cost of making root beer= 5/7= 0.71 pizzas

Because Becky has lower opportunity cost of making root beer, Becky has a comparative advantage in making root beer.

24) when country moves from A to B opportunity cost of producing 1 shirt= (20-15)/(10-0)= 0.5 bushels of corn.

When the country moves from C to D , opportunity cost of producing 1 shirt= (10-5)/(30-20)= 0.5 bushels of corn.